OpenStudy (anonymous):

A 2.0-gram bullet is shot into a tree stump. It enters at a speed of 3.00 x 104 cm/s and comes to rest after having penetrated 0.05 m in to the stump. What was the average force during the impact?

OpenStudy (anonymous):

@TheSmartOne ?

TheSmartOne (thesmartone):

\[\large\sf Force = \frac{ Mass \times (Final ~Velocity - Initial ~Velocity) }{ Change~in~time }\]

OpenStudy (anonymous):

force=unknown; mass = 2.0 grams; final velocity = 0; initial velocity = 3.00x104cm/2; change in time = 0.05 ?

TheSmartOne (thesmartone):

So they gave us that \(\sf Mass = 2.0 grams\\ Initial~Velocity = 3.00 \times 10^4 cm/s \\ Final~Velocity = 0\)

TheSmartOne (thesmartone):

I'm stumped on the time part... Hmmm

TheSmartOne (thesmartone):

Idk if we are using the correct formula either :/

OpenStudy (anonymous):

the struggle is real...

TheSmartOne (thesmartone):

This might help you I think :P https://answers.yahoo.com/question/index?qid=20150128075918AAoAJrs

OpenStudy (anonymous):

do you know anyone that might know how to help me solve this?

TheSmartOne (thesmartone):

You can maybe make a new post and I'll tag some users. They always come to my tag faster than another user since they know me. Haha